OpenStudy (tamara7301):

can someone help me with math plz :\ Which statements are true? Choose all answers that are correct. 8√=42√ 75−−√=53√ 45−−√=35√ 28−−√=47√

OpenStudy (alivejeremy):

u got any guess :D

OpenStudy (tamara7301):

not really i tried to simplify all of them but they arent the same but i think im simplifing wrong

OpenStudy (alivejeremy):

So how many can u pick

OpenStudy (tamara7301):

more then one

OpenStudy (alivejeremy):

8√=42√,45−−√=35√

OpenStudy (alivejeremy):

i would go with those :D

OpenStudy (alivejeremy):

O i did it wrong

hartnn (hartnn):

the screenshot helps! can you show your simplifications?

OpenStudy (tamara7301):

how did u do it:

OpenStudy (alivejeremy):

U had write it wrong lol

OpenStudy (tamara7301):

lel yea sorry

OpenStudy (tamara7301):

and hartann idk how

hartnn (hartnn):

try to factor 8?

OpenStudy (tamara7301):

2.82842712474619

OpenStudy (tamara7301):

do u mean sqr?

hartnn (hartnn):

i meant \(8 = 2\times 2\times2\)

OpenStudy (alivejeremy):

i think A is wrong

OpenStudy (tamara7301):

o

hartnn (hartnn):

similarly, can you factor 75, 45 and 28?

OpenStudy (tamara7301):

isnt b wrong?

OpenStudy (tamara7301):

i got 81 for b:\

OpenStudy (alivejeremy):

Be is right

OpenStudy (alivejeremy):

B*

OpenStudy (tamara7301):

how?im so confused

OpenStudy (alivejeremy):

Ask hartnn

OpenStudy (alivejeremy):

he can explain if its right or wrong

hartnn (hartnn):

let me do one for you, try doing others similarly \(\Large \sqrt{75} = \sqrt{3\times 5\times 5} = \sqrt 3 \times \sqrt{5\times 5} \\ \Large = \sqrt 3\times \sqrt{5^2} = \sqrt 3 \times 5 = 5\sqrt 3 \implies true\)

OpenStudy (tamara7301):

o ok

hartnn (hartnn):

try to understand the steps, and let me know if you don't get any of the steps

OpenStudy (tamara7301):

hows there 2 5's?

hartnn (hartnn):

75 = 3*25 = 3*5*5

OpenStudy (tamara7301):

ooo

OpenStudy (alivejeremy):

https://mathway.com/

OpenStudy (alivejeremy):

:D ^^^

hartnn (hartnn):

thats why the first step is to factor now try to factor 45 and 28

OpenStudy (tamara7301):

15*3=45 5*9=45

OpenStudy (tamara7301):

thx jay

OpenStudy (tamara7301):

by 3*5

hartnn (hartnn):

so we have \(\sqrt {45} = \sqrt{5} \times \sqrt 9\) and whats the square root of 9 ?

OpenStudy (tamara7301):

3

hartnn (hartnn):

yeah so we have \(\sqrt{45} = 3\sqrt 5 \implies true\)

hartnn (hartnn):

how about factoring 28?

OpenStudy (tamara7301):

2*14=28 4*7=28

hartnn (hartnn):

good so we have \(\sqrt{28} = \sqrt 4 \sqrt 7 \) and whats the square root of 4 ?

OpenStudy (tamara7301):

2

hartnn (hartnn):

\(\sqrt{28} = 2\sqrt 7\) so \(\sqrt{28}= 4\sqrt 7 \: \: is \: \: false.\)

OpenStudy (tamara7301):

o ok

hartnn (hartnn):

can you try with \(\sqrt 8\) ? hint 8 = 2*2*2

OpenStudy (tamara7301):

2*4=8 so its false??

hartnn (hartnn):

\(\sqrt 8 = \sqrt{4 \times 2} = 2 \sqrt 2\) so yea false

OpenStudy (tamara7301):

o ok